Presentation on theme: "Thursday/Friday 6-7 ECS scripts test in Orsay (Alberto,Vincent, Frederic and me) –automatic configuration (scripts are generated."— Presentation transcript:
Christophe.Suire@ipno.in2p3.fr Thursday/Friday 6-7 ECS scripts test in Orsay (Alberto,Vincent, Frederic and me) –automatic configuration (scripts are generated depending on the EquipmentId/DDL selected in DATE) –test with quadrant 9 –check results with mood ECS scripts tested up to 0-supp data taking ! equipmentId selected in DATE is correctly propagated configuration files (init,ldMarc,zSup etc… ) are still hand-made calibration script ready to be tested at cern (index was missing)
Christophe.Suire@ipno.in2p3.fr Sunday 9 Stand alone test/running –check St1 DDL (all 4) –check trigger sequence l2a and and l2ar –a few try with ecs script –check St2 DDL or Ch3Right (need LC2 on the corresponding LDC) nightly access to cavern : open ? Access will be needed during the week : 4 boards have to be replaced (~ 1 or 2 hour work, including tests) replacement of FRT2 in 2560 (pb with buses 17 and 31) buses 46,47 and 48 disconnected in 2562 After fixes all the St1 is there no power on C20 crate…. not possible to send trigger!!
Christophe.Suire@ipno.in2p3.fr Monday 10 Stand alone running –readout 1 DDL of station and test the dedicated ECS scripts –readout of the remaining DDLs –tests of the global ECS scripts parametrized by the DDL present in the equipmentID list no night shift nightly access to cavern : probable problem with trigger words (modify crocus code) ECS scripts ready to be tested DCS is not seen by DCS DIM = aldaqecs.cern.ch MCH DCS is broken SCR meeting : HMPID in global and SPD ready soon. L3 doors not closing (Wednesday ?) but dipole is being tested.
Christophe.Suire@ipno.in2p3.fr Tuesday 11 Stand alone running –possible to read chamber 1 with l2ar.seq but CDH has to be disabled –stand alone run with ECS ok –no trigger on CH2R ; most probably will need to change the cable DCA : fake state machine to make MCH available CDH problem originates from the setup of the L0 parameter (cable transmission only or FPGA crash) checking CDH demands a particular trigger sequence !! SCR meeting : L3 doors will close sometime… Dipole resistance lower than expected (factor 5 !!). DAQ dim server almost solved but still disconnected from DCS. HMPID+MTRG+SDD have run in global partition (TRD,TOF and FMD on the way). ACCORDE not ready.
Christophe.Suire@ipno.in2p3.fr Wednesday 12 1.5 10 6 events with CDH ok… load the proper trigger sequence sod.seq,l2a.seq,eod.seq tested along with CDH DCS control OK for DCA control zero-suppression almost ok and calibrations run looks fine St2 (ddl 2564) boot init and ldMarc ok (all manus respond) manu HS on BP 29 of CH1R First Run with Chamber 1 in global partition : smooth integration SCR meeting : L3 doors closing requires moving cable trays january; for now test L3 with ~100 gauss. Dipole resistance still too low (R > 1MΩ to reach nominal field), current raised up to 500A. DIM server problem fixed. SDD + TOF have run together (SDD+HMPID run fails). DCS fire alarms are back. TPC black events cannot be transferred to CASTOR (check limit size). ACCORDE has still trigger words issues. DAQ stop at 7am tomorrow to include monitoring machines in DB. Friday 3 to 7pm dedicated to busy measurements.
Christophe.Suire@ipno.in2p3.fr Thursday 13 Need to test the shuttle ; test new mood version prepare full st1 for 0-supp and calib then include in global access to test CH2R and add a FTD board in trigger crate. shuttle partially tested clearly a problem with CH2R trigger cable ; tried to pull a old one, didnt work. problems when inserting a new FTD board in trigger dispatching mood tested; OK. Log on aldaqdqm05-6-7-8-9 SCR meeting : too many people in CR (34 max), ~100 sometimes. L3 doors closing: January. Still low insulation resistance for the dipole. DAQ global run with 5 detectors. CTP not in yet. V0 and MUTRG inputs connected to CTP. [DCS integration not done yet for muon_trk : see boccioli/ivan ]. TPC rising up voltages to try to see particles. Progress for ACCORDE but not in global yet. OFFLINE started data replication to TIER1 (FZK,ccin2p3); ~50% of estimated p+p data size moved.
Christophe.Suire@ipno.in2p3.fr Friday 14 check St1 HV : tested up to 1650V nothing above 10 nA check FRT of CH1R (to be done this w/e) Run in global partition with 2 LDCs chamber 1 + chamber 3 right Measurements of busy and trigger rates in global partition : results need to be analysed (summary 40kHZ trigger, data rate 2kHz with ~400 ) pedestal run started from global reads local run control settings (be careful when adding/removing LDC). Each LDC must have its own pedestal/electronic calib config in DATE. SCR meeting : Use electronics logbook! L3 doors same stuff. Dipole 380 KΩ measured (apparently normal ~ sum of the « slices » meseaured independently). DAQ: probleme to rootify big TPC events (360MB). 6 detectors in global and ACCORDE entered global. DAQ/DCS will retry to have common DIM server. Idea to display the busy of detectors on a big screen. TPC HV on, pb with laser. TRD in global.
Christophe.Suire@ipno.in2p3.fr Saturday 15 Improving St1 readout (changing manu) Lot of problems with trigger (never can be sure where it comes from…) Solved problem with Ch2R : problem with a switch on the CRT check carefully the switches config local vs dispatching of the CRT 4 DDL in one LDC - Full Station 1- read chamber by chamber problem with ECS timeout in LOAD_FERO with 4 DDL detector not in the right state and the physic run cannot start some test runs (see runs list) SCR meeting : not intended….
Christophe.Suire@ipno.in2p3.fr Sunday 16 Taking data in global partition…. Offline group should take a look at the data. A few standalone runs to test the shuttle a MANU died during a run this has the consequence to mess up the loading of the thresholds. Vincent will improve robustness of scripts. MTRK in global (Ch2 with HT) runs very well run full station 1, 1 LDC fully loaded. Some timeout still to be fix. SCR meeting : CTP readout in global OK, will add big screen display of detector in busy state (Feb). HLT can read data flow from detector (including MTR and MCH), some ESD from TPC events (not analyzed yet). ITS in global, half of SDP, 21 ladders of SSD. TPC HV up, still a problem with events. MCH, I did a little bit of advertising /showing off since we reached all our goals, some congratulations ;-). MTR HV almost reached. ACORDE + small scintillator (below TPC) has some timing problem, bit pattern might not be readable in offline. HPMID problem with trigger info (?). Offline, pb with LCG data replication stopped unexpectedly. Future plan (tuesday) All detectors (except TPC) in DAQ test. If OK, add the TPC for run with DATE check of headers.
Christophe.Suire@ipno.in2p3.fr Monday 17 Dipole tests; will try to take pedestal runs to measure the influence on the noise if R. Tieulent is around, will try to measure noise when GMS is on. Dipole has been powered (up to 4000A) during standalone runs : no effects seen Pedestal runs with GMS : no visible effects on noise. multiple boot of CROCUS without (off/on) MUST be implemented Some stability problems on LV for St1 : must check current limit used at Orsay. SCR meeting : Dipole current up to 4000A. DAQ, large event problem was a root I/O issue (now buffering TPC data sector by sector). HLT read TPC event, signal visible with a ped+6 ADC cut; should be lowered. SPD all c-side powered, some really bad spikes in T° probably related to water pressure drop (bubbles ?). TPC reach gain ~10000. ACORDE, no coincidences with small scintillator (give up). MCH, I had to go in the FASS plug –back- trigger cable on CH2R… why was it unplugged ???
Christophe.Suire@ipno.in2p3.fr Tuesday 18 our machine aldaqacr25 crashed badly yesterday, took some time to recover (might be related to a memory pb) Magnet stable a 4000A, a few pedestal runs. Current increased at 5000A. Survey to measure displacement. Runs in global : rather stable Franco told me that timeout for COMPUTE_THRESHOLD script (maybe others) is now 60 s. Will need to discuss time and #events Correlated hit seen in run 14511 (see run list) SCR meeting : discuss tonight plan to satisfy DAQ commissioning and data taking. Run DAQ with the max of detector till 12 with acorde trigger. At midnight, ITS takes over with small scintillator trigger. TPC runs standalone with acorde trigger to accumulate data.
Christophe.Suire@ipno.in2p3.fr Wednesday 19 NO Work Leaving (Vincent and Christophe) Staying for the morning (Fred and Phillippe) SCR meeting :
Christophe.Suire@ipno.in2p3.fr Run list in Global Partition Run list must be visible on the alice-logbook https://alice-logbook.cern.ch/ (no access outside cern network; you may use remote desktop connection cernts.cern.ch ) –11445 : SSD, SDD, T0 & MCH. ~2500 events @ 25 Hz with 0-sup, full chamber 1 run pedestal #12362 –12367 : MCH + CTP. ~77 events @ 25 Hz with 0-sup, full chamber 1 + ½ chamber 3 (right) –12368 : MCH. ~3500(?) events @ 70 Hz with 0-sup, full chamber 1 + ½ chamber 3 (right) –a set of run made for trigger rate/busy have been recorded (see log book) –13022 : MTRK (chamber 2 physics HT@1600) + CTP, accorde trigger 1 HZ (coincidence) 471 events –13137-48-50-60 : SSD, MTRK (chamber 2 physics HT@1600), TRD [+ HLT] –HT varies between 1500 and 1600 test HT shuttle
Christophe.Suire@ipno.in2p3.fr Run list in Global Partition –13995 (along with 7 det) pulser trigger, HV@1600. Ch1+Ch2, ended with a big crash of our machine in acr.. –12490 (along with 4 det) ACORDE trigger, HV@1600. Ch1+Ch2, Idipole=4000A –14493 (along with 6 det) ACORDE trigger, 82471 events, HV@1600. Ch1+Ch2, no field –14511 (along with 6 det) ACORDE trigger, 11209 events, HV@1600. Ch1+Ch2, no field correlated hit seen… should be analysed associated pedestal run : 14452
Christophe.Suire@ipno.in2p3.fr Run 14511 : Hit in chamber 2
Christophe.Suire@ipno.in2p3.fr Run list (not in global) Local test Runs (pulser trigger @100 Hz) –12963 pedestal run HT@1000V –12964 run with voltage oscillation 1600->1500->1600->1500 20h24 -> 20h45 ~120kevents –12967-68-70-71-72-73-74-75-76-77 calibration runs (few errors problem with the end of data not received after 10s) HT@1600V –12979 pedestal run HT@1000V –12983 pedestal run HT@1000V –13719 pedestal run (reference for no GMS noise ) Ch1R, Idipole=1500A –13722 pedestal run 1000 event Ch1R, Idipole=1500A –13723 pedestal run 1000 event Ch1R, Idipole=1500A –13726 pedestal run 1000 event, GMS Running Ch1R, Idipole=1500A –13730 pedestal run 1000 event, GMS Running Ch1R, Idipole=1500A –13732 pedestal run 1000 event, Ch1R, Idipole=1500A –14259 pedestal run, magnet @4000A –14261 pedestal run, magnet @4000A –14263 pedestal run, magnet @4000A